OpenStudy (blondisaurus):

How would you got about simplifying this fraction raised to an exponent. (In comment section, will award medal)

OpenStudy (blondisaurus):

\[-\left( \frac{ 8 }{ 5 } \right)^{-2}\]

OpenStudy (usukidoll):

hmmm negative exponents aren't allowed... anything with a negative exponent is rewritten as \[x^{-2} =\frac{1}{x^2}\]

OpenStudy (usukidoll):

\[\frac{1}{(\frac{8}{5})^2}\]

OpenStudy (blondisaurus):

Ohh i forgot that rule

OpenStudy (usukidoll):

I'm ignoring the negative sign for now, but \[\frac{1}{\frac{64}{25}}\] and flip the second fraction

OpenStudy (usukidoll):

\[-\frac{25}{64}\]

OpenStudy (usukidoll):

now I put the - back on...

OpenStudy (blondisaurus):

I see! You're awesome. Thanks a bunch!

OpenStudy (usukidoll):

\[- \frac{1}{\frac{8}{5}}\] that 8/5 is squared ^^

OpenStudy (usukidoll):

my latex is driving me to the fritz machine today xD

OpenStudy (usukidoll):

\[-\frac{1}{\frac{64}{25}} \rightarrow -\frac{25}{64}\]

OpenStudy (blondisaurus):

Oh ok